# Conditional Enzyme Generation Using Protein Language Models with Adapters

### v1

@@ 4 Results
Table 2: Summary of datasets and splits used for training and evaluation of ProCALM. Unless otherwise noted, models were trained on the Swissprot Train set.
@@
Train Common ECs Generation Evaluation Randomly sampled ECs from the pool of ECs that correspond to >500 sequences in Swissprot Train n/a 24
Train Rare ECs Generation Evaluation Randomly sampled ECs from the pool of ECs that correspond to <10 sequences in Swissprot Train n/a 24
@@
4.1 ProCALM matches the performance of existing methods
Parameter efficient finetuning is advantageous due to its low computational cost and greater robustness to overfitting (Sledzieski et al., 2024), compared to full parameter training. Impressively, our Uniref-9B and Swissprot-1.5B models only required 240 and 40 A100-hours to train, respectively, whereas ZymCTRL required 15,000 A100-hours (Fig. 2D). By comparison, ZymCTRL trained all model parameters for approximately 50 billion tokens and required a higher memory footprint than our parameter-efficient training. Additionally, in ProCALM, held-out sequences demonstrate similar perplexities to the sequences associated with the most common ECs in the train set, which suggests that our model is not overfitting (Fig. 2C). By contrast, in ZymCTRL and our own experiments where we trained all model weights (Fig. A.3), the held-out sequences had significantly higher perplexity than common sequences in the train set.
Figure 2: ProCALM matches existing methods at generating sequences from common EC classes, at a lower cost. (A) The generation quality of ProCALM is similar to existing methods (i.e., ZymCTRL), as measured by the fraction of generated sequences that resemble a known enzyme (Valid Enzymes) and by the fraction of those enzymes with the correct EC conditioning (Correct EC). (B) Generated enzyme sequences have good diversity, as measured by the maximum sequence identity to a reference database (Mean Max Seq Identity) and the fraction of unique clusters at 90% sequence identity (90% clusters). (C) The perplexities of protein sequences from common EC classes in the train set are similar to those held-out at 90% and 70% identity, which suggests robustness to overfitting. (D) Number of GPU hours required to train ProCALM models is significantly lower than ZymCTRL. Swissprot-1.5B and Uniref-9B are ProCALM models trained on the Swissprot Train and Uniref datasets to 1.5 and 9 billion tokens, respectively, using the OH encoding of the EC number as a conditioning representation. Shorter training on the Swissprot dataset seems sufficient to achieve good performance. Error bars indicate standard deviation among 24 Train Common ECs.
4.2 ProCALM accommodates multiple types of conditioning
Figure 4: ProCALM can perform out-of-domain generalization for rare and unseen functions (EC classes). (A) Three different representations of enzyme function explored in this work. OH encodes the hierarchy of EC numbers as a one-hot encoding, DRFP encodes a fingerprint of reactions associated to an EC number, and CREEP is a learned representation that encodes multiple modalities related to the EC number. (B) Our models generate a non-negligible fraction of Valid Enzymes for rare ECs in the train set and ECs that are entirely held out from the train set. (C) While almost none of the generated sequences using held-out ECs as conditioning mapped exactly to the target conditioning, (D) the generated sequences look more similar to the target EC. Accuracy level refers to the correctness of an EC. For example, if the target was 1.1.1.1, an EC of 1.2.1.2 would have an accuracy level of 1 and an EC of 1.1.1.2 would have an accuracy level of 3. (E) Perplexities are consistent for protein sequences from various splits, which suggests that the models are robust to overfitting. Models were trained to 1.5 billion tokens. Error bars indicate standard deviation among 24 ECs.
ProCALM can use other types of information, such as taxonomy, to condition sequence generation. We trained two additional models, one conditioned on taxonomy, and one conditioned jointly on taxonomy and EC. For the jointly conditioned model, we used a modified architecture, where parallel adapters accept multiple sources of conditioning information (Fig. A.2). In an ablation study, we found that parallel adapters work better than merging the joint conditioning in the conditioning encoder and using a shared adapter (Fig. A.6).
Overall, the taxonomy conditioned and jointly conditioned models enriched generation for the desired condition(s) (Fig. 3), showing that ProCALM is easily adaptable. Performance is highest for bacterial sequences, as they constitute the majority of the training data (Fig. A.1). However, learning multiple (EC and taxonomy) conditions sacrifices performance slightly compared to models trained to learn either EC or taxonomy conditioning alone.
4.3 ProCALM generalizes to out-of-distribution conditions
Generative models able to generate sequences for unseen functions would be highly impactful. For example, enzyme engineering often involves finding enzymes that can perform interesting non-native activities (Yang et al., 2024a; b; Chen & Arnold, 2020). We explored ProCALM’s ability to generate sequences conditioned on rare and held-out EC numbers; the latter task has not been explored before. This setup reflects a real-world scenario where certain enzyme functions have not been annotated, but their corresponding genes have been sequenced. We also explored ProCALM’s flexibility to accommodate different representations of enzyme function. Namely, we encoded the EC number as a OH encoding of the EC hierarchy, a continuous fingerprint representation of reactions associated with an EC number (DRFP) (Probst et al., 2022), and a multimodal embedding from contrastive learning (CREEP) (Yang et al., 2024b) – visualized in Fig. 4A with more details in Section A.1.
Overall, ProCALM can generate a substantial fraction of valid enzymes under this challenging conditioning (Fig. 4B). While a negligible number of generated sequences mapped to the correct EC number completely, the functions associated with the generated sequences were generally more similar to the target function (Fig. 4C-D). Using more continuous (DRFP and CREEP) representations of enzyme function did not result in better generalization performance. Held-out EC numbers with more similar functions to the training set generally yielded sequences with more similar functions to the target conditioning (Fig. A.7), and in general, the CREEP representation seems to capture somewhat different information than OH and DRFP. Finally, the perplexities of sequences associated with rare and held-out ECs are comparable to the train set, suggesting that our models are not prone to overfitting (Fig. 4E). Still, there is signifcant room for future research and improvement here.
Generating unseen combinations of taxonomy and enzymatic activity would also be interesting for applications such as discovering gene-editing enzymes in new organisms (Burstein et al., 2016). Thus, we next used ProCALM to generate bacterial sequences from several common EC classes, where all bacterial sequences in those EC numbers were held-out during model training. This scenario reflects a hypothetical real-world situation where a eukaryotic enzyme has not been discovered and annotated in a prokaryote. For certain EC classes, ProCALM could generate bacterial sequences (Fig. 5). Generation quality is better when the model is prompted with bacteria as conditioning (as opposed to other taxa), suggesting that the correct generations are not due to chance (Fig. A.8). Overall, ProCALM seems to be able to learn the distribution of bacterial sequences, allowing for some level of OOD generalization.
5 Discussion
We have shown that conditional adapters are an efficient, flexible, and generalizable approach for enabling language models to perform steerable sequence generation. Our specific implementation, ProCALM, demonstrates strong performance for enzyme sequence generation conditioned on known EC numbers, at a significantly lower computational cost than prior approaches. ProCALM easily adapts to different types of conditioning, such as enzyme function and taxonomy, and to different representations for the same condition, such as OH, DRFP, and CREEP representations of enzyme function. Our approach is also parameter efficient, resulting in lower memory usage and potentially faster training.
Most usefully, ProCALM can perform previously unexplored generation tasks, such as generating from unseen EC classes. Still, performance is limited, and there is significant room for improvement in this area. For certain EC classes, ProCALM was able to generate bacterial sequences corresponding to the correct EC, despite not having bacterial sequences belonging to those EC classes in the train set. 7.1.1.2 and 7.1.1.9 are transmembrane enzymes part of large complexes, but it is not clear why these particular functions were successfully generated. In the future, a more in depth examination of success and failure modes for ProCALM will be interesting.
For known functions, it is relatively straightforward to use function prediction tools such as BLAST to filter generated sequences to the valid ones with correctly conditioned function, though future work will benefit from better function prediction oracles for evaluation (Yu et al., 2023; Ayres et al., 2024; Huo et al., 2024). Future evaluation will also include more detailed analysis of generation quality (Ye et al., 2024) using various tools including structure prediction (Johnson et al., 2024; Alamdari et al., 2023; Nicolini et al., 2024).
In general, we believe that the most interesting research direction for future improvement is generation of sequences with OOD functions and/or properties. In real-world application, evaluation is particularly intractable for these unannotated functions, so it will be even more critical to increase the hit rate for enzyme discovery on those functions. Having a smooth and navigable latent space for OOD functions will be especially important, along with training data encompassing more comprehensive annotations of natural and engineered proteins. While we explored various representations here based on fingerprints and contrastive learning, further exploration of multiple modalities such as structure, sequence, function, and text could lead to breakthroughs (Yang et al., 2024b; Mikhael et al., 2024; Ramesh et al., 2022; Yin et al., 2024). For convenience, we finetuned our model using annotated proteins in Swissprot, but greater sequence diversity by incorporating sequences from unannotated and metagenomic datasets could be beneficial in the future.
Overall, we have demonstrated that using conditional adapters to finetune language models offers several advantages, including computational efficiency, flexibility to many types of conditioning, and generalization for OOD generation. This approach could easily be applied to other autoregressive PLMs (Hesslow et al., 2022; Ferruz et al., 2022). ProCALM is publicly available, and we encourage researchers to apply our approach to other language models and further explore its potential for conditional protein generation.
Acknowledgments
The authors thank Jacopo Teneggi for preliminary explorations that laid the foundation for this work, as well as Stephen Nayfach, Joel Beazer, and Sarthak Jain for helpful discussions.
Appendix A Appendix
A.1 Datasets and Representations
@@
Uniref and Swissprot were downloaded from Uniprot (The UniProt Consortium et al., 2023) as all sequences with EC numbers on June 17, 2024. We left Uniref unprocessed for training. For Swissprot, we only considered the sequences corresponding to EC numbers with associated reactions from CARE (Yang et al., 2024b), and we first resampled it by weighting each sample with $\frac{1}{1+\ln{(n)}}$ relative probability of being sampled, where $n$ is the size of the cluster that a sequence belongs to when clustered at 50% identity using MMseqs2 (Steinegger & Söding, 2017). Effectively, this up-samples the sequences from smaller clusters. Afterward, we generated the held-out splits outlined in Table 2 and Fig. A.1. The remaining non held-out sequences in Swissprot Train were then used for training most models. For Fig. 5, we updated Swissprot Train slightly by also holding out bacterial sequences associated with the 9 EC classes explored.
@@
Different representations of EC numbers were obtained for the Swissprot dataset. For OH, each of the four levels of the EC number were OH-encoded and concatenated, resulting in a 630-dimensional vector. The EC number is a hierarchical scheme consisting of four levels where each subsequent level describes enzyme function more specifically. For DRFP, DRFP representations (Probst et al., 2022) with dimenion 2048 were obtained for all reactions associated to each EC based on CARE (Yang et al., 2024b), and the EC was encoded as the mean DRFP representation. DRFP is a fingerprint associated with the set change from reactants to products and has shown state-of-the-art ability for reaction classification. For CREEP, we used the model trained with default parameters on the medium train-test split for Task 2, which was trained using contrastive learning to align representations of protein sequences, reactions, and textual description (Yang et al., 2024b). From this model, we extracted representations for the textual description associated to each EC number.
@@ A.2 Architecture and Training Details
We finetuned ProGen2-base, which is a decoder-only autoregressive language model consisting of 27 layers, an embedding dimension of 1536, and 764 million parameters (Nijkamp et al., 2023). By default, the conditioning encoder consisted of two layers with a dimension of 256 each. The output was then concatenated with a low-rank projection of the language hidden state (dimension 16). Each adapter MLP consisted of 3 layers with dimension 2*(16 + 256) and a final layer with dimension (256 + 16). For the default ProCALM model, the adapter layers and conditioning encoder together consisted of approximately 70 million additional parameters, depending on the condition dimension. The adapter parameters dominated compared to the conditioning encoder weights. During parameter efficient training, the original ProGen2 weights were fixed, and the adapter and conditioning encoder were trained (pseudocode provided by Ruffolo et al. (2024a)). During full finetuning, all weights were trained. For the Small model, dimensions were reduced by half and only two layers were used in the adapter MLP. For the parallel adapter architecture, conditioning information was passed through separate adapters before being summed as the embedding update (Fig. A.2), resulting in approximately twice the number of trainable parameters. Future work could explore the impact of scaling to larger ProGen2 models.
@@ A.3 Evaluation Metrics
Sequences were generated with probabilistic decoding with a top-p value of 0.95 and a temperature of 0.3. To determine the validity and function of generated sequences, we used Diamond BLASTp (Buchfink et al., 2021). Our reference database consisted of 550k proteins from Swissprot, not just enzymes. For the main text figures, 900 sequences were generated for each unique condition, and for the remaining figures, 225 sequences were generated for each condition to get good statistics. Valid Enzymes were defined as those mapping to a hit in the reference database with default parameters, with at least 80% alignment coverage. A valid enzyme was determined to have a Correct EC if its BLASTp hit in the reference database mapped to the target EC number.

### v2

@@ A.1 Datasets and Representations
Table A.1: Summary of datasets and splits used for training and evaluation of ProCALM on EC-related tasks. Unless otherwise noted, models were trained on the Swissprot Train set using the ProGen2-Base model (764 million parameters). For the text-guided generation task, the same training dataset as ProteinDT was used, and no dataset resampling or splitting was performed.
@@
Train Common ECs Generation Evaluation Randomly sampled ECs from the pool of ECs that correspond to ¿500 sequences in Swissprot Train n/a 24
Train Rare ECs Generation Evaluation Randomly sampled ECs from the pool of ECs that correspond to ¡10 sequences in Swissprot Train n/a 24
@@
Uniref and Swissprot were downloaded from Uniprot (The UniProt Consortium et al., 2023) as all sequences with EC numbers on June 17, 2024. We left Uniref unprocessed for training. For Swissprot, we only considered the sequences corresponding to EC numbers with associated reactions from CARE (Yang et al., 2024b), and we first resampled it by weighting each sample with $\frac{1}{1+\ln{(n)}}$ relative probability of being sampled, where $n$ is the size of the cluster that a sequence belongs to when clustered at 50% identity using MMseqs2 (Steinegger & Söding, 2017). Effectively, this up-samples the sequences from smaller clusters. Afterward, we generated the held-out splits outlined in Table A.1 and Fig. A.1. The remaining non held-out sequences in Swissprot Train were then used for training most models. For Fig. 6, we updated Swissprot Train slightly by also holding out bacterial sequences associated with the 9 EC classes explored.
@@
Different representations of EC numbers were obtained for the Swissprot dataset. For OH, each of the four levels of the EC number were OH-encoded and concatenated, resulting in a 630-dimensional vector. The EC number is a hierarchical scheme consisting of four levels where each subsequent level describes enzyme function more specifically. For DRFP, DRFP representations (Probst et al., 2022) with dimension 2048 were obtained for all reactions associated to each EC based on CARE (Yang et al., 2024b), and the EC was encoded as the mean DRFP representation. DRFP is a fingerprint associated with the set change from reactants to products and has shown state-of-the-art ability for reaction classification. For CREEP, we used the model trained with default parameters on the medium train-test split for Task 2, which was trained using contrastive learning to align representations of protein sequences, reactions, and textual description (Yang et al., 2024b). We note that there is no data leakage in the CREEP representation, as it was trained with the same held-out EC numbers. From this model, we extracted representations for the textual description associated to each EC number.
@@ A.2 Architecture and Training Details
Unless otherwise noted, we finetuned ProGen2-Base, which is a decoder-only autoregressive language model consisting of 27 layers, an embedding dimension of 1536, and 764 million parameters (Nijkamp et al., 2023). By default, the conditioning encoder consisted of two layers with a dimension of 256 each. The output was then concatenated with a low-rank projection of the language hidden state (dimension 16). Each adapter MLP consisted of 3 layers with dimension 2*(16 + 256) and a final layer with dimension (256 + 16). For the default ProCALM model, the adapter layers and conditioning encoder together consisted of approximately 70 million additional parameters, depending on the condition dimension. The adapter parameters dominated compared to the conditioning encoder weights. During parameter efficient training, the original ProGen2 weights were fixed, and the adapter and conditioning encoder were trained (pseudocode provided by Ruffolo et al. (2024a)). During full finetuning, all weights were trained. For the Small model, dimensions were reduced by half and only two layers were used in the adapter MLP. For the parallel adapter architecture, conditioning information was passed through separate adapters before being summed as the embedding update (Fig. A.2), resulting in approximately twice the number of trainable parameters. We also briefly explored the impact of scaling to larger ProGen2 models.
@@ A.3 Evaluation Metrics
Sequences were generated with probabilistic decoding with a top-p value of 0.95 and a temperature of 0.3, except for the text-guided generation task, which used a temperature of 1.0. To determine the validity and function of generated sequences, we used Diamond BLASTp (Buchfink et al., 2021). Our reference database consisted of 550k proteins from Swissprot, not just enzymes. For the main text figures, 900 sequences were generated for each unique condition, and for the remaining figures (including the text-guided conditioning task), 225 sequences were generated for each condition to get good statistics. Valid proteins were defined as those mapping to a hit in the reference database with default parameters, with at least 80% alignment coverage. A valid protein was determined to have a Correct EC or function if its BLASTp hit in the reference database mapped to the target EC number or natural language prompt.
@@ A.4 Additional Results
Figure A.5: Losses and perplexities of ProCALM models, for various model sizes and training and test splits. Overall, the ProCALM model based on ProGen2 XLarge shows lower losses and perplexities. All models were trained using Swissprot Train and default parameters. Perplexity was measured on the models trained to 1.5 billion tokens.
Figure A.6: Various metrics used to evaluate the quality, diversity, and novelty of generated sequences, as a function of model size. 3 different pretrained models were used: ProGen2-Base (764 million parameters), ProGen2-Large (2.7 billion parameters), and ProGen2-XLarge (6.4 billion parameters). Overall, scaling does not seem to have significant effects on the quality of generated sequences.
@@
Table A.2: Example prompts used to evaluate natural language-conditioned generation from PLMs.
Common Prompts
carbamoyl phosphate + L-aspartate = H(+) + N-carbamoyl-L-aspartate + phosphate Pyrimidine metabolism; UMP biosynthesis via de novo pathway; (S)-dihydroorotate from bicarbonate: step 2/3. Belongs to the aspartate/ornithine carbamoyltransferase superfamily. ATCase family.
ATP + L-cysteine + tRNA(Cys) = AMP + diphosphate + L-cysteinyl-tRNA(Cys) Binds 1 zinc ion per subunit. Monomer. Belongs to the class-I aminoacyl-tRNA synthetase family.
ATP-dependent specificity component of the Clp protease. It directs the protease to specific substrates. Can perform chaperone functions in the absence of ClpP. Component of the ClpX-ClpP complex. Forms a hexameric ring that, in the presence of ATP, binds to fourteen ClpP subunits assembled into a disk-like structure with a central cavity, resembling the structure of eukaryotic proteasomes. Belongs to the ClpX chaperone family.
Involved in protein export. Acts as a chaperone by maintaining the newly synthesized protein in an open conformation. Functions as a peptidyl-prolyl cis-trans isomerase. [protein]-peptidylproline (omega=180) = [protein]-peptidylproline (omega=0) About half TF is bound to the ribosome near the polypeptide exit tunnel while the other half is free in the cytoplasm. Consists of 3 domains; the N-terminus binds the ribosome, the middle domain has PPIase activity, while the C-terminus has intrinsic chaperone activity on its own. Belongs to the FKBP-type PPIase family. Tig subfamily.
Binds directly to 16S ribosomal RNA. Belongs to the bacterial ribosomal protein bS20 family.
ATP + L-histidine + tRNA(His) = AMP + diphosphate + H(+) + L-histidyl-tRNA(His) Homodimer. Belongs to the class-II aminoacyl-tRNA synthetase family.
Rare Prompts
Involved in the regulation of gene expression by abscisic acid, stress factors and by components of stress signal transduction pathways. Transcription factor that binds to the GCC-box pathogenesis-related promoter element. Part of a transcriptional repressor complex including a histone deacetylase. Interacts with SIN3 and HDA19. Contains a slightly degenerated ERF-associated amphiphilic repression (EAR) motif, which may be involved in the activity of transcriptional repression. Phosphorylated by PKS3. Belongs to the AP2/ERF transcription factor family. ERF subfamily.
Functions as a component of the Five Friends of Methylated CHTOP (5FMC) complex; the 5FMC complex is recruited to ZNF148 by methylated CHTOP, leading to desumoylation of ZNF148 and subsequent transactivation of ZNF148 target genes (By similarity). Component of the PELP1 complex involved in the nucleolar steps of 28S rRNA maturation and the subsequent nucleoplasmic transit of the pre-60S ribosomal subunit (By similarity). May play a role during development (By similarity). Component of the 5FMC complex, at least composed of PELP1, LAS1L, TEX10, WDR18 and SENP3; the complex interacts with methylated CHTOP and ZNF148. Interacts with NOL9. Component of the PELP1 complex, composed of at least PELP1, TEX10 and WDR18. The complex interacts with pre-60S ribosome particles. Mainly found in the nucleoplasm, with low levels detected in the cytoplasmic and chromatin fractions. Belongs to the WD repeat IPI3/WDR18 family.
Involved in the final reduction of the elongation cycle of fatty acid synthesis (FAS II). Catalyzes the NADH-dependent reduction of a carbon-carbon double bond in an enoyl moiety that is covalently linked to an acyl carrier protein (ACP). It can use both crotonyl-CoA and crotonyl-ACP. a 2,3-saturated acyl-[ACP] + NAD(+) = a (2E)-enoyl-[ACP] + H(+) + NADH a 2,3-saturated acyl-CoA + NAD(+) = a (2E)-enoyl-CoA + H(+) + NADH (2E)-butenoyl-[ACP] + H(+) + NADH = butanoyl-[ACP] + NAD(+) butanoyl-CoA + NAD(+) = (2E)-butenoyl-CoA + H(+) + NADH Weakly inhibited by triclosan. Lipid metabolism; fatty acid biosynthesis. Monomer. Belongs to the TER reductase family.
Translation initiation regulator which represses repeat-associated non-AUG (RAN) initiated translation probably by acting as a competitive inhibitor of eukaryotic translation initiation factor 5 (EIF5) function (PubMed:29470543, PubMed:34260931). Enhances histone H4 gene transcription but does not seem to bind DNA directly (PubMed:11524015). Expressed in day 3 embryo. Belongs to the BZW family.